OpenStudy (anonymous):

the diameter of a circle is 24. find the area in terms of pi

OpenStudy (anonymous):

(12)^2pi

jimthompson5910 (jim_thompson5910):

Diameter: d = 24 Radius: r = d/2 = 24/2 = 12 Area: A = pi*r^2 A = pi*12^2 A = pi*144 A = 144pi So the exact area is 144 pi square units.